OpenStudy (anonymous):

can you guys plz help simplify each expression d^-4/e^-7,and 6m^-8n/p^-1 ( i'll give out a medal)

OpenStudy (anonymous):

\[\frac{ d^{-4} }{ e^{-7} }=\frac{ 1 }{ d^4 }*\frac{ e^7 }{ 1 }=\frac{ e^7 }{ d^4 }\]\[\frac{ 6m^{-8}n }{ p^{-1} }=\frac{ 6n }{ m^8 }*p^1=\frac{ 6np }{ m^8 }\] I don't know if 'e' is a variable or the euler's number so I left it like that.

OpenStudy (anonymous):

@wild

OpenStudy (anonymous):

e is the is a variable

OpenStudy (anonymous):

Ok so that's fine then.

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

i dont get how you did the last problem

OpenStudy (anonymous):

never mind i got